# Generalized Models for Inflationary Preheating: Oscillations and Symmetries

### v2

@@ B.3 Subdominant growth from second quartic coupling
The trajectories in $(A,q)$ space for the second quartic coupling tend to be farther in the stable regions compared to trajectories for the trilinear case with the same $(A_{0},q_{0})$, resulting in subdominant growth. The trajectories and growth rates are shown in Figures 10 and 11. Notably, growth rates from the trilinear and second quartic interactions can be comparable when $V(\phi)\not\supset\lambda\phi^{4}/4$ depending on choices of the $q$ parameters in each Dufaux et al. (2006); trilinear growth remains dominant with our choice of setting ${q_{\sigma}}_{\mathrm{max}}\approx{q_{g_{2}}}_{\mathrm{max}}$.
Figure 10: Trajectory examples in $(A,q)$ space for $m_{\chi}/\omega_{0}=0$ (top row) and $m_{\chi}/\omega_{0}=2$ (bottom row) for second quartic preheating. The black $A=2q$ line roughly separates the unstable region above from the stable region below. The solid lines correspond to the mixed case $m\neq 0,\ \lambda\neq 0$, the dashed lines to purely quartic $m=0,\ \lambda\neq 0$, and the dotted lines to purely massive $m=0,\ \lambda\neq 0$. The three columns correspond from left to right to $u_{0}\equiv\lambda\Phi_{0}^{2}/m^{2}=1,10,1000$ in the mixed case, and the parameters for the massive and quartic cases are picked to match $(A_{0},q_{0})$ of the mixed cases. The colors correspond to different wavenumbers $k_{0}/\omega_{0}$, listed in Figure 11. Notice that for $m_{\chi}=0$, the pure quartic $(A,q)=(A_{0},q_{0})=$constant because conformal symmetry is preserved.
Figure 11: Growth rates $\gamma$ along trajectories as a function of scale factor $a$ corresponding to the trajectories in Figure 10. The solid lines correspond to the mixed case $m\neq 0,\ \lambda\neq 0$, the dashed lines to purely quartic $m=0,\ \lambda\neq 0$, and the dotted lines to purely massive $m=0,\ \lambda\neq 0$. The three columns correspond from left to right to $\lambda\Phi_{0}^{2}/m^{2}=1,10,1000$ in the mixed potential, and the parameters for the massive and quartic cases are picked to match $(A_{0},q_{0})$ of the mixed cases. The colors correspond to different wavenumbers $k_{0}/\omega_{0}$; note that for the massive case, $\omega_{0}=m$, for the quartic case, $\omega_{0}\approx 0.85\sqrt{\lambda}\Phi_{0}$, and for the mixed case, $\omega$ is the function shown in equation (54). These growth rates are lower than those from trilinear preheating in Figure 5, even with similar or slightly greater $q_{\mathrm{max}}$.
B.4 Explosive growth from third quartic coupling
@@
where the $q$-parameter increases by $q\to q_{\mathrm{eff}}=q\chi_{k}(t+T)/\chi_{k}(t)\approx e^{\gamma}q$ every cycle, so the degree of exponential instability itself increases exponentially. This is desirable for efficient preheating. However, these extreme growth rate increases necessitate the inclusion of backreaction and rescattering effects to correctly calculate, which is outside the scope of and a reasonable next step to this paper.
@@ Acknowledgements
We thank Nemanja Kaloper for useful conversations. S.W. thanks the Simons Center for hospitality. This research was supported in part by DOE grant DE-FG02-85ER40237. A.B. was partially supported by NSF grant DMS-2103026, and AFOSR grants FA 9550-22-1-0215 and FA 9550-23-1-0400. L.B. and F.C.A. also thank the Leinweber Center for Theoretical Physics at the University of Michigan. We also thank the referee for useful suggestions on improving the first draft.

### v3

@@ C.3 Subdominant growth from second quartic coupling
Figure 15: Trajectory examples in $(A,q)$ space for $m_{\chi}/\omega_{0}=0$ (top row) and $m_{\chi}/\omega_{0}=2$ (bottom row) for second quartic preheating. The black $A=2q$ line roughly separates the unstable region above from the stable region below. The solid lines correspond to the mixed case $m\neq 0,\ \lambda\neq 0$, the dashed lines to purely quartic $m=0,\ \lambda\neq 0$, and the dotted lines to purely massive $m=0,\ \lambda\neq 0$. The three columns correspond from left to right to $u_{0}\equiv\lambda\Phi_{0}^{2}/m^{2}=1,10,1000$ in the mixed case, and the parameters for the massive and quartic cases are picked to match $(A_{0},q_{0})$ of the mixed cases. The colors correspond to different wavenumbers $k_{0}/\omega_{0}$, listed in Figure 16.
[t]
Figure 16: Growth rates $\gamma$ per oscillation time along trajectories as a function of scale factor $a$ corresponding to the trajectories in Figure 15. The colors correspond to different wavenumbers $k_{0}/\omega_{0}$; note that for the massive case, $\omega_{0}=m$, for the quartic case, $\omega_{0}\approx 0.85\sqrt{\lambda}\Phi_{0}$, and for the mixed case, $\omega$ is the function shown in equation (55). These growth rates are lower than those from trilinear preheating in Figure 5, even with similar or slightly greater $q_{\mathrm{max}}$.
@@
The trajectories in $(A,q)$ space for the second quartic coupling tend to be farther in the stable regions compared to trajectories for the trilinear case with the same $(A_{0},q_{0})$, resulting in subdominant growth. The trajectories and growth rates are shown in Figures 15 and 16. Notably, growth rates from the trilinear and second quartic interactions can be comparable when $V(\phi)\not\supset\lambda\phi^{4}/4$ depending on choices of the $q$ parameters in each Dufaux et al. (2006b); trilinear growth remains dominant with our choice of setting ${q_{\sigma}}_{\mathrm{max}}\approx{q_{g_{2}}}_{\mathrm{max}}$.
Notice that for $m=m_{\chi}=0$, the pure quartic potential scenario, $(A,q)=(A_{0},q_{0})=$ constant because conformal symmetry is mostly preserved, as was shown in Greene et al. (1997). Following the idea in Appendix B, we note that while $\gamma$ is constant for this potential, $\gamma/H$ decreases with time, another reason for the favorable efficiency of the trilinear interaction, for which we showed $\gamma/H$ increases with time.
C.4 Explosive growth from third quartic coupling
@@
where the $q$-parameter increases by $q\to q_{\mathrm{eff}}=q\chi_{k}(t+T)/\chi_{k}(t)\approx e^{\gamma}q$ every cycle, so the degree of exponential instability itself increases exponentially. This is desirable for efficient preheating. However, these extreme growth rate increases necessitate the inclusion of backreaction and rescattering effects to correctly calculate, which is outside the scope of and a reasonable next step from this paper.
@@ Acknowledgements
We thank Nemanja Kaloper for useful conversations. S.W. and J.T.G. thank the Simons Center for hospitality. This research was supported in part by DOE grant DE-FG02-85ER40237. A.B. was partially supported by NSF grant DMS-2103026, and AFOSR grants FA 9550-22-1-0215 and FA 9550-23-1-0400. L.B. and F.C.A. also thank the Leinweber Center for Theoretical Physics at the University of Michigan. J.T.G. is supported in part by the National Science Foundation, PHY-2309919. We also thank the referee for useful suggestions on improving the first draft. We’d like to thank our reviewers for constructive comments that improved our thoroughness and quality.